# Good News for Polynomial Root-finding

### v35

@@ 1.4 Some Benefits of Black Box Polynomial Root-finders
(i) Black box polynomial root-finders are particularly efficient for $d$th degree polynomials $p$ that can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (ops), such as Mandelbrot-like polynomials $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, and the sums of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, Horner’s algorithm uses $2d$ ops to evaluate at a point $x$ a general polynomial $p=p(x)$ of (1.1) given with its $d+1$ coefficients, and this bound is optimal (see [85, 66] or [135, Section “Pan’s method"]). One can evaluate $p$ at $d_{+}>d$ points and then interpolate to it, but this would destroy sparsity, blow up the precision and bit operation cost of the computations, and dramatically slow down root-finding in a region containing a small number of roots.
(ii) The known efficient polynomial root-finders extensively shift (translate) and scale variable $x$ and in particular routinely apply map (1.4) to reduce root-finding in any disc $D(c,\rho)$ to root-finding in the unit disc $D(0,1)$. All this, however, is done at the price of losing sparsity and blowing up computational precision and bit operation complexity of root-finding.
(iii) The same problems severely limit application of the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (see [49, 82, 109] for the history), which compute polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$, as follows:
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i+1}}(x^{2}):=(-1)^{d}p_{2^{i}}(x)p_{2^{i}}(-x),~{}i=0,1,\dots$ (1.5)
Among various helpful properties of iterations (1.5), root-squaring also squares the distances from the origin to the roots, thus moving the external roots away from the unit circle $C(0,1)$, that is, strengthens its isolation from the roots.
By avoiding computation of the coefficients of $t(x)$ and $p_{2^{i}}(x)$ of maps (1.4) and (1.5), respectively, black box polynomial root-finders keep all benefits of using these maps but avoid the cited notorious problems.
For yet another benefit we can apply black box root-finders to polynomials in Bernstein form [71], whose evaluation is known to be numerically stable.
Remark 1.2.
Polynomial root-finding in the 19-th and well into the 20-th century relied on these iterations combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.4) from any complex point (see [82, 109, 111] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests in [111] with standard test polynomials have consistently demonstrated efficiency of a single particular implementation of this approach, but even the straightforward implementation of iterations (1.5) promises to be highly efficient as well.
@@ 1.5 Our complexity estimates
We adopt the model of [73] but otherwise rely on distinct approach and techniques (with few exceptions listed below). Indeed, the root-finder of [73] cannot help solve even Problem 1: even in the explicit setting, where the coefficients of $p$ are given, repeatedly deflating $p$, by peeling off large roots, would accumulate errors that grow too quickly, and it is not entirely clear how to even do this in the black box setting. Even if this recipe worked it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [89, 24, 31]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [92, 99] for general polynomial root-finding. [^8] For root-finding problems one should assume that $b$ exceeds $d$ because the precision of representation of $p$ (input length) increases by a factor of $d$ in routine shifts and scaling of the variable $x$.
Recursive combination of [73] with the Inverse Power Iterations [46] would enable approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors. One can be challenged to estimate the increased computational complexity of the solution of this extended problem of matrix computations, but we study other approaches.
In particular we accelerate approximation of $m$ zeros of $p$ for $m\leq d$ by a factor of $m$ by means of incorporation of fast multipoint polynomial evaluation of [40, 81, 65, 78], which we could not use if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation.
As in [73] we reduce root-finding to the evaluation of $p(x)$ at sufficiently many points. The ops involved in this evaluation are more numerous than the other ops involved in our algorithms, even in the case where $p(x)=x^{d}$.
Unlike [73] we use Las Vegas randomization – to control with a high probability (whp) precision and bit-operation complexity of our computation. In particular we readily prove that whp $p(x)$ and $p^{\prime}(x)$ are not too close to 0 for a random variable $x$, and so we perform division safely whp when we approximate Newton’s ratio NR$(x):=\frac{p(x)}{p^{\prime}(x)}$ or its reciprocal – Newtons Inverse Ratio NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at random points $x$.
For these basic steps of our root-finders we fix a complex $\delta$ whose absolute value is sufficiently small in context and approximate $p^{\prime}(x)$ with the ratio $\frac{p(x)-p(x-\delta)}{\delta}$. This ratio is the first order divided difference, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$ by virtue of Taylor-Lagrange’s formula, [^9] Unlike [73] we do not need to approximate higher order derivatives and thus involve no higher order divided differences; by using Las Vegas randomization, however, we still ensure whp the same bound $O(b\log(d))$ of [73] on the computational precision. and so $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$. Substitute the ratio for $p^{\prime}(x)$ and obtain that
Our root-finders approximate the roots lying in the unit disc $D(0,1)$ together with some roots lying nearby. At the end we only keep the roots lying in the disc $D(0,1+\epsilon)$ but at the intermediate steps may approximate some extraneous roots in a larger domain, but we modify subdivision in Sec. 1.6.2 so that all extraneous roots lie in the disc $D(0,1+\beta)$ for any fixed positive $\beta$.
@@
Hereafter $m=m(\beta)$ denotes the number of the zeros of $p$ lying in the disc $D(0,1+\beta)$ for a fixed positive constant $\beta$ of our choice. [^10] See our Cor. 1.2 and compare more restrictive choice of $\beta=2$ in [22, 23].
Our Las Vegas randomized root-finders support the Main Theorems 1-4 below.
@@
(Main Theorem 1.) One can solve Problems 1 and $*{}^{*}$ by applying Las Vegas randomized queries of the polynomial evaluation oracle with a computational precision $O(b\log(d))$ at $\tilde{O}(dm)$ locations for $b:=\log_{2}(1/\epsilon)$ and $m$ of Def. 1.1.
By following [73] one can readily extend Main Theorem 1 to matrix eigenvalue approximation:
@@
(Main Theorem 2.) One can solve Problems 2 and $*{}^{*}$ at Las Vegas randomized bit operation cost
$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m)$ (1.7)
for $m$ of Def. 1.1, $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM),  [^11] The current records are about 2.7734 for exponents of feasible MM [86, 103, 105, 69, 54], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 35]. The exponents below 2.77 are only reached for unfeasible MM – supported by algorithms that are slower than the classical $d\times d$ MM unless $\log_{2}(\log_{2}(\log_{2}(d))))>100$, say. and $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$ for the Frobenius norm $||M||_{F}$ of a matrix $M$.
Thm. 1.1 implies the bound $\tilde{O}(bd^{2}m)$ on Las Vegas randomized bit operation complexity of Problem 1 because one can evaluate any polynomial $p$ of degree $d$ in $2d$ ops and perform an op with the precision of $s$ bits by using $O(s\log(s)\log(\log(s)))=\tilde{O}(s)$ bit operations, which extends the celebrated bound
$\mu(s):=O(s\log(s)\log(\log(s)))=\tilde{O}(s)$ (1.8)
on the number of bit operations for multiplication of two integers modulo $2^{s}$ [137] (see [66, 130, 131] for this extension).
In the case of the Mandelbrot polynomials $p:=M_{k}(x)$ and the sums of a constant number of shifted monomials one can immediately decrease by a factor of $d/\log(d)$ both of the above upper bounds of $2d$ ops and $\tilde{O}(bd^{2}m)$ bit-operations.
Even in the worst case of general polynomial $p$ given with its coefficients, however, we decrease the upper bound on the bit operation cost of our root-finders by roughly a factor of $d$ – to match the near optimal bound of [92, 99] in our next theorem. We achieve this by combining our root-finders supporting Main Theorem 1 with the known algorithms for fast multipoint polynomial evaluation [40, 81, 65, 78].
@@
(Main Theorem 3.) One can solve Problems 0 and $*{}^{*}$ at Las Vegas randomized bit operation cost $\tilde{O}((b+d)dm)$ for $b:=\log_{2}(1/\epsilon)$, $m$ of Def. 1.1, and general polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
Finally, we decrease our upper bounds of Thms. 1.1-1.3 by a factor of $d$ under an additional assumption on the disposition of the zeros of $p$ and eigenvalues of $M$, namely, under the following model: [^12] Its assumption is a little stronger than that of our Def. B.1, supporting such a decrease in [92, 99], but it is in good accordance with the results of extensive tests in [57, 60] for standard test polynomials (see Sec. 13.4.1 and the beginning of Sec. 4.1).
Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.
@@
In view of Remarks 1.1 and C.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
1.6 Our algorithms: design and analysis (outline)
For a polynomial $p$ given in monomial basis (with its coefficients) our deterministic algorithms of [92, 99] support the Main Theorems 3 and 4, but our present Las Vegas randomized algorithms does this in the case of a black box polynomial as well.
1.6.1 Subdivision root-finders and reduction to e/i tests
We depart from the popular subdivision root-finders, traced back to [144, 50, 124, 97], which extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction. Let a square contain precisely $m>0$ zeros of $p$ (with multiplicities). Call it suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test, to decide whether the square contains any root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a).
@@ 1.6.1 Subdivision root-finders and reduction to e/i tests
Observation 1.1.
A subdivision iteration applied to a square with $m$ zeros of $p$ (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every zero of $p$ within the distance of one half of their diameter.
@@
Readily verify claim (i). Claim (ii) holds because a zero of $p$ can lie in at most four suspect squares (and even in at most two unless it is the shared vertex of the four). Claim (iii) holds because a subdivision iteration never discards a square containing a root. ∎
@@
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ (and diameter $\Delta\sqrt{2}$) containing precisely $m$ zeros of $p$. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ zeros of $p$ lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
1.6.2 Soft e/i tests and compression of a Region of Interest
Instead of testing whether a suspect square $S$ contains some zeros of $p$, Henrici in [50] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. We adopt his approach, which implies that the zeros of $p$ lying outside the square but in the disc can affect the result of our e/i test. Furthermore, in the presence of rounding errors no e/i test can decide whether a zero of $p$ lies in or outside a disc if it actually lies on or very close to its boundary circle. To address both problems we fix $\sigma>1$ and apply soft e/i tests as follows.
@@ 1.6.2 Soft e/i tests and compression of a Region of Interest
Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion. Exclusion certifies that the square $S$ contains no zeros of $p$ and therefore should be discarded from subdivision process. $\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step. Criteria for both exclusion and $\sigma$-soft inclusion can hold simultaneously (see Fig. 1b), but the test ends as soon as it verifies any of them.
Let us estimate how transition to $\sigma$-soft e/i tests affects Cor. 1.1.131313We counter the adverse impact of rounding errors by increasing $\sigma$ but only need tiny increase for small $\epsilon$ and will ignore it in our estimates.
Observation 1.2.
(i) The number $n(\sigma)$ of the $\sigma$-covers of suspect squares that contains a fixed zero of $p$ is $O(1)$, that is, bounded by a constant, if so is $\sigma$. (ii) In particular $n(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
Remark 1.3.
We only use $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$ in our root-finders.
Seeking roots in the unit disc $D(0,1)$ we apply subdivision iterations to its minimal covering square $S(0,1)=\{x:~{}|\Re(x)|\leq 1,~{}|\Im(x)|\leq 1\}$, having side length 2 (see Fig. 1b) and $\sigma$-cover $D(0,\sigma\sqrt{2})$.
Observation 1.3.
Let $U_{i}=U_{i}(\sigma)$ denote the union of the $\sigma$-covers of all suspect squares processed at the $i$-th subdivision step with $\sigma$-soft e/i tests. Then (i) $U_{i+1}(\sigma)\subset U_{i}(\sigma)$, that is, domain $U_{i}(\sigma)$ shrinks as $i$ increases, (ii) $U_{i}(\sigma)\subset D(0,2)$ for all $i\geq 0$ if $\sigma<\sqrt{2}$, and (iii) $U_{h}(\sigma)\subset D(1+\beta)$ for any fixed positive $\beta$ and some $h=h(\sigma,\beta)$, which is a constant, $O(1)$, if so is $\sigma$.
Under the Random Root Model the numbers of roots lying in the discs $D(0,1)$ and $D(0,2)$ differ by a constant factor $f=O(1)$ whp, and so for $\sigma<\sqrt{2}$ the estimates of Cor. 1.1 hold up to that factor whp, due to claim (ii) of the latter observation.
Based on its claims (i) and (iii) we next extend Cor. 1.1 deterministically, by incorporating the following two recipes of Extension and Delay for discarding squares in subdivision process.
(a) Extend discarding: In addition to discarding squares according to $\sigma$-soft e/i tests, discard all squares that lie outside the disc $D(0,1)$.
(b) Delay discarding: For a fixed positive $\beta$ begin applying $\sigma$-soft e/i tests and discarding squares only at the $h$th subdivision step for $h=h(\sigma,\beta)$ of claim (iii) of Observation 1.3.
Clearly, the $h$th subdivision iteration may process up to $4^{h}=O(1)$ suspect squares, although a little fewer for larger $h$ if the extension of discarding is applied.
Now we can extend Cor. 1.1 as follows.
@@
Fix $\beta>0$ and $\sigma>1$, let the disc $D(0,1+\beta)$ contain precisely $m=m(\beta)$ zeros of $p$ (cf. Def. 1.1), and apply subdivision iterations (with $\sigma$-soft e/i tests under the recipes of extension and delay of discarding) to a minimal square $S(0,1)$ covering the unit disc $D(0,1)$. Then (i) the $k$-th iteration involves at most $n(\sigma)m$ suspect squares for $n(\sigma)$ of Observation 1.2, $n(\sigma)\leq 4$ for $\sigma<\sqrt{2}$; the centers of those squares approximate within $2^{1.5-k}$ all the zeros of $p$ lying in the disc $D(0,1)$ and (ii) the first $k$ subdivision iterations involve no suspect squares and $\sigma$-soft e/i tests for $k<h=h(\sigma,\beta)$ and at most $(k-h-1)n(\sigma)m+4^{h}$ suspect squares and $\sigma$-soft e/i tests for $k\geq h$.
Remark 1.4.
We extend Cor. 1.1 for subdivision process applied to the unit disc $D(0,1)$ but can similarly do this for any fixed Region of Interest.
1.6.3 Cauchy-based root-counting and e/i tests
We have highly efficient heuristic algorithms based on the DLG iterations (1.5) that closely approximate the smallest distance to the roots from a complex point (see Remark 1.2); this can be immediately extended to soft e/i tests.
Next we cover efficient e/i tests that also enable formal support of our Main Theorems 1– 4.
Let #$\#(\mathcal{D})$ denote the number of zeros of $p$ lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that

### v36

@@ 1.4.1 Why we should use them
(iii) Black box polynomial root-finders are particularly efficient where a $d$th degree polynomial $p$ can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (hereafter (ops) such as a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [89, 69] or [139, Section “Pan’s method"]).
1.4.2 What we adopt and what cannot adopt for black box polynomial root-finding
The known polynomial root-finders, except [76], extensively use the coefficients of $p$, and we must be careful in adopting their techniques for Problems 1 and 1.
On the other hand, the root-finder of [76] cannot help solve even Problem 1. Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors (notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say). This is true even if one performs deflation in the black box setting implicitly, based on Eqn. (3.3) of Sec. 3. Moreover, even in that setting the coefficient swell in deflation is translated into swelling of the values of the high order derivatives involved in the algorithm of [76].
Even if the algorithm of [76] worked for Problem 1 it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [93, 25, 33]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [96, 103] for general polynomial root-finding.
Recursive combination of [76] with the Inverse Power Iterations [48] could lead to approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors. One can be challenged to estimate the increased complexity of this extended computations, but we study other approaches.
By applying a root-finder to Given a polynomials $p$ of (1.1) given with its coefficients We accelerate known algorithms for Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation algorithms of [42, 84, 68, 81], but we could not use them if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation. Thus, instead of applying high order Newton’s iterations of [76] and deflation, we recall and extend subdivision iterations of [149, 52, 128, 101].
@@ 1.5 Our complexity estimates
$m=m(\mathbb{D},\beta)$ for $\beta>0$ is the number of roots of polynomial $p(x)$ in the $\beta$-neighborhood of a domain $\mathbb{D}$ on the complex plane.
Main Theorems 1-4 below hold for $m$ of Def. 1.1, any fixed constant $\beta>0$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise.
@@
(Main Theorem 1.) One can solve Problems 1 and 1 by querying the polynomial evaluation oracle at $\tilde{O}(m^{2})$ locations with a computational precision $O(b\log(d))$.
@@
(Main Theorem 2.) One can solve Problems 2 and 2 at Las Vegas randomized bit operation cost
$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m^{2})$ (1.6)
for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM). [^4] The current records are about 2.7734 for exponents of feasible MM [90, 107, 109, 72, 57], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 37]. The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
@@
(Main Theorem 3.) One can solve Problems 0 and 0 at Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.555(i) The assumption of the model is a little stronger than that of our Def. 24.1, supporting such a decrease in [96, 103]. (ii) Random coefficient models may seem more natural for polynomial root-finding but are useless because no known probability distribution withstands shift of the variable $x$.
@@
In view of Remarks 1.1 and 25.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
2 Accelerated subdivision polynomial root-finders: overview and reduction to soft exclusion/inclusion tests
For a polynomial $p$ given in monomial basis (with its coefficients) the deterministic algorithm of [96, 103] supports our Main Theorems 3 and 4, but not our Main Theorems 1 and 2. This is because the algorithm of [96, 103] cannot be applied to a black box polynomial, unlike our present Las Vegas randomized algorithms, which we outline next and elaborate upon in Parts II–IV. Namely, in the next two sections we reduce our root-finders to recursive application of exclusion/inclusion (e/i) test, which decides whether a fixed complex disc contains roots or not. In Secs. 3.3.2 and 3.4 we outline our two approaches to devising e/i tests for a black box polynomial and unit disc $\mathcal{D}=D(0,1)$. We cover these two approaches in some detail in Part II, made up of Secs. 612, and Part III, made up of Secs. 1318, respectively, and describe yet another fast e/i test in Sec. 23.4. Our root-finders of Parts II and III incorporate these e/i tests and are quite fast but still only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 1923 and outlined in Sec. 4, we fix both deficiencies, thus proving our Main Theorems. In Sec. 5.4 we list some tentative directions to further progress.
2.1 Subdivision root-finders and reduction to e/i tests
We depart from the popular subdivision root-finders, traced back to [149, 52, 128, 101]. They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction.
Next we outline them.
(A) First let a square on the complex plane contain precisely $m>0$ roots (with multiplicities) and let us approximate them. Call the square suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test in order to decide whether the square contains a root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a). In Sec. 9 and then in our Part IV we use a natural extension of e/i tests to $m$-tests, which decide whether a disc in the complex plane covering at most $m$ roots does or does not contain $m$ roots.
@@ 2.1 Subdivision root-finders and reduction to e/i tests
Observation 2.1.
A subdivision iteration applied to a square with $m$ roots (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
@@
Readily verify claim (i). Claim (ii) holds because a root can lie in at most four suspect squares, and even in at most two unless it is the shared vertex of the four. Claim (iii) holds because a subdivision iteration never discards a square containing a root. ∎
@@
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta\sqrt{2}$ containing precisely $m$ roots. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ roots lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
(B) Seeking roots in any fixed domain $\mathbb{D}$ in the complex plane, we apply subdivision iterations to its minimal covering square $S$ and while performing them also discard all squares that do not intersect the domain $\mathbb{D}$, even if they contain roots.
Then the $(\Delta_{k}\sqrt{2})$-neighborhood of $\mathbb{D}$ covers the union $\mathbb{U}_{k}$ of all suspect squares defined after the $k$th subdivision iteration, while $\Delta_{k}=\Delta_{0}/2^{k}$ fast converges to 0 as $k$ increases. Hence the neighborhood fast converges to the domain $\mathbb{D}$: the width of the neighborhood decreases by a factor of $2^{k}$ in $k$ subdivision iterations, still at the price of performing at most $4km$ e/i tests.
Remark 2.1.
[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection as follows: (i) recursively partition (subdivide) every suspect rectangle into two congruent rectangles, to each apply e/i test, and then discards a rectangle if it contains no roots but call it suspect and subdivides it otherwise, (ii) stop where the rectangles have diameter at most $2\epsilon$, so that their centers approximate the roots within $\epsilon$.
2.2 Subdivision iterations with soft e/i tests
Instead of testing whether a suspect square $S$ contains some roots, Henrici in [52] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle. To address both problems we fix $\sigma>1$ and apply soft e/i tests.
@@ 2.2 Subdivision iterations with soft e/i tests
[See the end of Sec. 6 for a generalization, covering in particular soft $m$-tests.] Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion. Exclusion certifies that the square $S$ contains no roots and therefore should be discarded from subdivision process. $\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step. Criteria for both exclusion and $\sigma$-soft inclusion can hold simultaneously (see Fig. 1b), but the test ends as soon as it verifies any of them.
Observation 2.2.
Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root. Then $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$, and $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
@@
Cor. 2.1 still holds if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
3 Fast soft e/i tests
3.1 Background: Newton’s Inverse Ratio and its approximation
We reduce e/i tests of subdivision root-finders to the multipoint evaluation of the ratio
${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}},$ (3.1)
which we call Newton’s inverse ratio. To prove the latter well-known and much used equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
The arithmetic operations (hereafter ops) involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
By following [76] we fix a complex $\delta$, with $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula. Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and
We deduce the same bound $O(b\log(d))$ of [76] on computational precision of our root-finders, but its derivation is greatly simplified versus [76] because we only use the first order derivatives and the first order divided differences.
Unlike [76] we estimate bit-operation cost of our root-finders and allow random choice of $x$ to control the value of the ratio $\frac{|p(x)-p(x-\delta)|}{p(x)\delta}$ with a high probability (whp). This is Las Vegas randomization, and so is any randomization in polynomial root-finding, as we show next.
3.2 Background: Error detection and correction
Proposition 3.1.
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
One can accelerate low precision numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) of [50, 28, 32].
3.3 Cauchy-based root-counting and e/i tests
3.3.1 Introductory comments
Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that